find the total payment,giventhe cost,tax rate,and tip rate for $51.80,6.7%tax,20%tip
Answer by user_dude2008(1862)

6.7% = 0.067 and 20% = 0.20

Total Cost = Cost + Cost*(Tax Rate) + Cost*(Tip Rate) = Cost * (1 + Tax Rate + Tip Rate)


Total Cost = Cost * (1 + Tax Rate + Tip Rate)

Total Cost = 51.80 * (1 + 0.067+ 0.20)

Total Cost = 51.80 * 1.267

Total Cost = 65.6306


Total cost is $65.63